2. Apple claims that the average wait time for a customer calling the Apple Care support line is 175 seconds. A simple random sample of 45 customers was selected, and their average wait time was found to be 187 seconds. Assume the population standard deviation of wait time is 50 seconds. Is there enough evidence to conclude that the average wait time for a customer calling the Apple Care support line is more than 175 seconds? Use α = 0.01.

a) State the null and alternative hypotheses in the context of the question.

b) Calculate the value of the test statistic. Check the required condition(s) to validate the test statistic.

c) Find the p-value.

d) What conclusion can we draw?
Answer #1

\\Given,\\ Level~of~significance(\alpha)=0.01\\\\ sample~Size(n)=45\\\\ Standard~Deviation(\sigma)=50\\\\ Sample~Mean(\bar{x})=187\\\\ .....................................................\\ a.)\\Null~Hypothesis~H_0:\mu=175\\\\ Alternative~Hypothesis~H_1:\mu>175\\\\ ........................................................\\ b.)Test~Statistic(Z)=\frac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\frac{187-175}{50/\sqrt{45}}=1.61\\\\ .........................................................\\ c.) P-value=P(|Z|>1.61)\\\\ ~~~~~~=2*P(Z<-1.61)\\\\=0.1074\\\\ Ans:0.1074\\\\ ............................................................\\\\ d.)as~p-value~is~greater~than significance~level(\alpha=0.01), Fail ~to~Reject~Null~Hypothesis.\\\\ Conclusion:Hence, ~there~is~not~Enough~evidence~to~support~the~claim~that~wait ~time ~for ~a ~customer ~calling ~the ~Apple ~Care ~support ~line~ is ~more~ than~ 175 ~seconds.